# A Simple Approach to Unifying Diffusion-based Conditional Generation

### v1

@@ 1 Introduction
As shown in Fig. 1, with the same model but different sampling schedules, UniCon can do: 1) controllable image generation in the form of ControlNet, Readout Guidance, and Loose Control, 2) estimation, and 3) joint generation. We train our models based on a large text-to-image model for several different modalities (depth, edges, human poses, image appearance) and show that the behavior of our single model is similar to or better than specialized methods using standard image quality and alignment metrics. We demonstrate significant improvements over prior unified models in conditional generation, training efficiency, and generation flexibility. We also show that UniCon can combine multiple models for multi-signal conditional generation or switch our model to other base model checkpoints. Our models are trained in about 13 hours on 2-4 Nvidia A800 80G GPUs, adding 15% parameters to the base model.
@@ 2 Related Work
Controllable Generation. Fine-tuning text-to-image diffusion models to conditional image generation on signals beyond text has gained significant popularity (Huang et al., 2023; Zhang et al., 2023b; Ye et al., 2023; Mou et al., 2023; Sohn et al., 2023). ControlNet (Zhang et al., 2023b) trains a control network attached to pre-trained diffusion models to incorporate condition signals, such as edge maps, segmentation masks, and pose estimation. Based on ControlNet, LooseControl (Bhat et al., 2024) generalizes depth conditioning to loose depth maps that specify scene boundaries and object positions. Readout-Guidance (Luo et al., 2024) proposes a new control scheme by adding prediction heads to internal features and guiding the generation by the predicted condition signal. Conditional editing tasks, such as DiffEdit  (Couairon et al., 2023), have enhanced conditional image manipulation by applying diffusion-based models for inpainting and editing tasks. Instead of one specific control behavior, our proposed framework provides diverse controlling abilities through different sampling strategies.
@@ 5 Results
We base our experiments on Stable Diffusion (Rombach et al., 2022) (SD), a large-scale text-to-image diffusion model. We train 5 UniCon models, Depth, SoftEdge, Human-Pose (Pose), Human-Identity (ID), and Appearance on SDv1-5. The first three pair an image with a spatially aligned condition image. Following Luo et al. (2024), we train Depth, SoftEdge models on 16k images from PascalVOC (Everingham et al., 2012) and Pose model on a subset with 9k human images. Depth, soft edge, and pose images are estimated by Depth-Anything-v2 (Yang et al., 2024), HED (Xie & Tu, 2015) and OpenPose (Cao et al., 2019). We train the ID model on 30k human face images from CelebA (Liu et al., 2015) and use images with the same identity as training image pairs. The Appearance model, aiming at transferring one image’s appearance to another, is trained on 6k videos from Panda70M (Chen et al., 2024b). We randomly select frames in the same video clip as training image pairs. Please refer to the Appendix for other training and inference details.
@@ 5.1 Main results
Fig 4 shows sample results from all UniCon models. Apart from common spatial aligned control signals, our ID and Appearance models work for loosely correlated images. With the ID model, we can generate images of the same person in the condition image and utilize the input prompt to specify appearance and style. The Appearance model is trained on video frames, aiming at generating images with a similar appearance. In practice, we can set the joint cross attention weight (multiply to $\mathbf{F}^{joint}$ in Eq. 4) to allow more image variations, e.g. the same wolf with a different pose.
Table 1: Quantitative conditional generation comparison. We compare Readout-Guidance (Luo et al., 2024), ControlNet (Zhang et al., 2023b), and UniCon(Ours) on 6K validation images. For a fair comparison, all methods are trained on PascalVOC with the same annotator.
@@
The conditional generation performance is related to the choice of training datasets and condition annotators. For a fair comparison, we train all models on the same dataset (PascalVOC) and annotator. We do not compare with JointNet because there is no training code provided. As shown in Tab. 1, our UniCon achieves similar or better performance than Readout-Guidance (Luo et al., 2024) and ControlNet (Zhang et al., 2023b) on FID and condition fidelity over all modalities.
We finetune our Depth model for 5K steps with Depth-Anything-V2-Metric (Yang et al., 2024) as the annotator for metric depth evaluation. As shown in Tab. 2, our Depth-Metric model performs similarly or better than MiDaS (Ranftl et al., 2020) and DPT (Ranftl et al., 2021). There is a margin between our model and Marigold (Ke et al., 2024), which fine-tunes the whole diffusion model and solely focuses on depth estimation. In comparison, our model trains fewer parameters and targets on unified conditional generation.
@@
Comparison with JointNet. We compare our method to the most relevant baseline, JointNet (Zhang et al., 2023a), which also supports multiple conditional generation tasks for image and depth. As demonstrated in Fig. 3, our method shows better results in image-to-depth, partial condition, and depth guidance, with extra abilities of rough depth condition. Quantitatively, UniCon achieves superior depth estimation results compared to JointNet (Tab. 2). Beyond the capabilities of JointNet, our method extends to handle non-pixel-aligned conditions (Fig. 4), supports multi-signal conditions (Fig. 8), and offers the flexibility of checkpoint switching (Fig. 8).
We also emphasize the differences in training methodology. JointNet trains on 65M images from the COYO-700M dataset (Byeon et al., 2022) on 64 NVidia A100 80G GPUs for 24 hours, including an additional fine-tuning stage for inpainting mask. It adds 889M parameters (100% of the base model SD2.1). In contrast, our UniCon-Depth trains on 16K images on 2 Nvidia A800 80G GPUs for 13 hours and adds 125M parameters (15% of the base model SDv1.5).
@@ Appendix B Training
Depth, SoftEdge, Pose. For spatially aligned conditions, we follow Readout-Guidance (Luo et al., 2024) to train on PascalVOC (Everingham et al., 2012). Depth and SoftEdge model is trained on 16K images and Pose model is trained on 9K images of humans. To obtain the condition input, we first annotate training images with existing estimation methods. Depth, soft edge, and pose images are estimated by Depth-Anything-v2 (Yang et al., 2024), HED (Xie & Tu, 2015) and OpenPose (Cao et al., 2019). We encode all estimated modalities as images, following the annotators in ControlNet (Zhang et al., 2023b). For SoftEdge, we follow ControlNet (Zhang et al., 2023b) to quantize the edge maps into several levels to remove possible hidden patterns. We train Depth, SoftEdge models for 20K steps with batch size 32 and Pose model for 10K steps. Training 20K steps costs about 13 hours on two NVIDIA A800 80G GPUs. To adapt our Depth model for metric depth estimation, we further fine-tune the model for 5K steps on images annotated with Depth-Anything-v2-Metric. The produced Depth-Metric model is used for the depth estimation evaluation in Tab. 2.
@@ D.2 Sampling Setting
For the qualitative comparison in Fig. 3, we generate specialized methods with their default sampling setting. JointNet (Zhang et al., 2023a) does not support guidance in their official implementation. Therefore, we adopt the same guidance scheme on their model to generate the depth guidance sample. We tune our sampling setting to generate each sample, such as the noise level added to the input depth for the rough depth condition task.
@@ Appendix E Additional Results
In Fig. 9, we show that we can interpolate the initial noise levels for input ${\mathbf{x}},{\mathbf{y}}$ to gradually change our sampling behavior in ${\mathbf{y}}$-to-${\mathbf{x}}$, rough ${\mathbf{y}}$-to-${\mathbf{x}}$, joint ${\mathbf{x}},{\mathbf{y}}$ generation, rough ${\mathbf{x}}$-to-${\mathbf{y}}$, and ${\mathbf{x}}$-to-${\mathbf{y}}$.

### v2

@@ 1 Introduction
As shown in Fig. 1, with the same model but different sampling schedules, UniCon can do: 1) controllable image generation in the form of ControlNet, Readout Guidance, and Loose Control, 2) estimation, and 3) joint generation. We train our models based on a large text-to-image model for several different modalities (depth, edges, human poses, identity) and show that the behavior of our single model is similar to or better than specialized methods using standard image quality and alignment metrics. We demonstrate significant improvements over prior unified models in conditional generation, training efficiency, and generation flexibility. We also show that UniCon can combine multiple models for multi-signal conditional generation or switch our model to other base model checkpoints. Our models are trained in about 13 hours on 2-4 Nvidia A800 80G GPUs, adding 15% parameters to the base model.
@@ 2 Related Work
Controllable Generation. Fine-tuning text-to-image diffusion models to conditional image generation on signals beyond text has gained significant popularity (Huang et al., 2023; Zhang et al., 2023b; Ye et al., 2023; Mou et al., 2023; Sohn et al., 2023). ControlNet (Zhang et al., 2023b) trains a control network attached to pre-trained diffusion models to incorporate condition signals, such as edge maps, segmentation masks, and pose estimation. Based on ControlNet, LooseControl (Bhat et al., 2024) generalizes depth conditioning to loose depth maps that specify scene boundaries and object positions. Readout-Guidance (Luo et al., 2024) proposes a new control scheme by adding prediction heads to internal features and guiding the generation by the predicted condition signal. DAG (Kim et al., 2022) guides a diffusion model to generate geometrically plausible images using depth prior. Conditional editing tasks, such as DiffEdit  (Couairon et al., 2023), have enhanced conditional image manipulation by applying diffusion-based models for inpainting and editing tasks. Instead of one specific control behavior, our proposed framework provides diverse controlling abilities through different sampling strategies.
@@ 5 Results
We base our experiments on Stable Diffusion (Rombach et al., 2022) (SD), a large-scale text-to-image diffusion model. We train 4 UniCon models, Depth, SoftEdge, Human-Pose (Pose), Human-Identity (ID) on SDv1-5. The first three pair an image with a spatially aligned condition image. Following Luo et al. (2024), we train Depth, SoftEdge models on 16k images from PascalVOC (Everingham et al., 2012) and Pose model on a subset with 9k human images. Depth, soft edge, and pose images are estimated by Depth-Anything-v2 (Yang et al., 2024b), HED (Xie & Tu, 2015) and OpenPose (Cao et al., 2019). We train the ID model on 30k human face images from CelebA (Liu et al., 2015) and use images with the same identity as training image pairs. In addition, we train an auxiliary Appearance model to cooperate with other models, as shown in Fig 9. It is trained on random frame pairs from 6k videos in Panda70M (Chen et al., 2024b), aiming at generating images with similar visual appearance.
@@ 5.1 Main results
Fig 4 shows sample results from all UniCon models. Apart from common spatial aligned control signals, our ID model works for loosely correlated human images. With the ID model, we can generate images of the same person in the condition image and utilize the input prompt to specify appearance and style.
Figure 5: User study of conditional generation performance. We compare our UniCon against Readout-Guidance (Luo et al., 2024), ControlNet (Zhang et al., 2023b) for spatially-aligned conditions (depth, softedge, pose) and against IP-Adapter-Face (Ye et al., 2023) for ID condition.
Table 1: Quantitative comparison of conditional generation for spatially-aligned conditions. We compare Readout-Guidance (Luo et al., 2024), ControlNet (Zhang et al., 2023b), and UniCon (Ours) on spatially-aligned conditions depth, softedge, and pose. For a fair comparison, all methods are trained on PascalVOC with the same annotator.
@@
We conduct a user study to compare UniCon models against their corresponding methods to evaluate the conditional generation performance. As shown in Fig. 5, UniCon demonstrates comparable performance to specialized methods in human preference (Fig. 5).
In addition, we compare UniCon for spatially-aligned conditions (depth, softedge, pose) against ControlNet (Zhang et al., 2023b), Readout-Guidance (Luo et al., 2024) in terms of FID and condition fidelity metrics. As shown in Tab. 1, our UniCon achieves similar or better performance than Readout-Guidance (Luo et al., 2024) and ControlNet (Zhang et al., 2023b) on FID and condition fidelity over all modalities.
We finetune our Depth model for 5K steps with Depth-Anything-V2-Metric (Yang et al., 2024b) as the annotator for metric depth evaluation. As shown in Tab. 2, our Depth-Metric model performs similarly or better than MiDaS (Ranftl et al., 2020) and DPT (Ranftl et al., 2021). There is a margin between our model and Marigold (Ke et al., 2024), which fine-tunes the whole diffusion model and solely focuses on depth estimation. In comparison, our model trains fewer parameters and targets on unified conditional generation.
@@
Comparison with JointNet. We compare our method to the most relevant baseline, JointNet (Zhang et al., 2023a), which also supports multiple conditional generation tasks for image and depth. As demonstrated in Fig. 3, our method shows better results in image-to-depth, partial condition, and depth guidance, with extra abilities of rough depth condition. Quantitatively, UniCon achieves superior depth estimation results compared to JointNet (Tab. 2).
@@ 6 Discussion
Acknowledgements. This work was supported in part by the National Natural Science Foundation of China (62322113, 62376156).
@@ Appendix B Training
Depth, SoftEdge, Pose. For spatially aligned conditions, we follow Readout-Guidance (Luo et al., 2024) to train on PascalVOC (Everingham et al., 2012). Depth and SoftEdge model is trained on 16K images and Pose model is trained on 9K images of humans. To obtain the condition input, we first annotate training images with existing estimation methods. Depth, soft edge, and pose images are estimated by Depth-Anything-v2 (Yang et al., 2024b), HED (Xie & Tu, 2015) and OpenPose (Cao et al., 2019). We encode all estimated modalities as images, following the annotators in ControlNet (Zhang et al., 2023b). For SoftEdge, we follow ControlNet (Zhang et al., 2023b) to quantize the edge maps into several levels to remove possible hidden patterns. We train Depth, SoftEdge models for 20K steps with batch size 32 and Pose model for 10K steps. Training 20K steps costs about 13 hours on two NVIDIA A800 80G GPUs. To adapt our Depth model for metric depth estimation, we further fine-tune the model for 5K steps on images annotated with Depth-Anything-v2-Metric. The produced Depth-Metric model is used for the depth estimation evaluation in Tab. 2.
@@ Appendix D Evaluation
Table 5: Relative depth estimation performance on DA-2K evaluation benchmark. We compare UniCon-Depth model against Marigold (Ke et al., 2024), Geowizard* (Fu et al., 2025), Depth Anything V1 (Yang et al., 2024a) (DA V1), ZoeDepth (Bhat et al., 2023), Depth Anything V2 (Yang et al., 2024b) (DA V2). *: results from Depth Anything V2.
Method Marigold* Geowizard* DAT V1* ZoeDepth UniCon-Depth (Ours) DAT V2 (ViT-L)*
Accuracy (%) 86.8 88.1 88.5 89.1 90.5 97.1
Table 6: Comparison between synchronous and asynchronous sampling scheduling. Synchronous: denoising image and depth together while replacing depth input with noisy GT depth at each step. Asynchronous (default): denoising image with clean depth input. We test on UniCon-Depth for depth-to-image generation.
Method FID CLIP Score AbsRel $\delta_{1}$
Sync 14.78 32.45 17.43 77.00
Async 13.21 32.11 9.26 91.02
Figure 10: Inpainting and editing samples. We show diverse inpainting and editing results using UniCon-Depth and UniCon-Edge.
@@ D.2 Sampling Setting
For the qualitative comparison in Fig. 3, we use the default sampling setting for specialized methods. JointNet (Zhang et al., 2023a) does not support guidance in their official implementation. Therefore, we adopt the same guidance scheme on their model to generate the depth guidance sample. We tune our sampling setting to generate each sample, such as the noise level added to the input depth for the rough depth condition task.
@@ Appendix E Additional Results
Table 7: Inference Speed Comparison for conditional generation. We compare UniCon against ControlNet (Zhang et al., 2023b) on denoising iterations per second.
Method ControlNet UniCon (Remove Joint Modules) UniCon
iteration/s 10.82 7.65 5.02
Table 8: Latency Comparison for Depth Estimation. We compare UniCon against Marigold (Ke et al., 2024), ZoeDepth (Bhat et al., 2023), and Depth Anything V2 (Yang et al., 2024b) on denoising iterations per second, where Marigold and UniCon are diffusion-based. The latency is tested on one A800 GPU.
Method Marigold Marigold(LCM) ZoeDepth Depth Anything V2 UniCon
latency (A800) 2.3s 372ms 289ms 91ms 4.0s
User Study. We conduct a user study on conditional generation performance for depth, softedge, pose, and identity conditions. For each model, users are asked to select their preferred image that aligns with the given image and text condition. We show the result in Fig. 5, where UniCon produces comparable results against other specialized methods.
For the Appearance model, evaluating it independently is challenging because (1) Applying it in isolation often produces outputs that closely resemble the input image. We must adjust the condition weight or text prompt for specific use cases. Its effectiveness is most apparent when combined with other models. Please check Fig. 1,4,9,11 for its use cases. (2) A suitable baseline for comparison is currently unavailable.
Synchronous and asynchronous sampling schedules. Our disentangled timestep sampling scheme during training enables flexible asynchronous sampling schedules at inference time, which is how UniCon handles diverse generation tasks. In Tab. 6, we compare the depth-to-image generation performance under synchronous and asynchronous sampling schedules. The results demonstrate that asynchronous noise level scheduling offers clear advantages in depth conditional generation.
For the synchronous setting, we denoise image and depth together while replacing depth input with noisy GT depth at each step. For the asynchronous setting, we denoise the image with clean depth input, which is only possible due to our training strategy.
Quantitative depth estimation. Though our UniCon model still lags behind state-of-the-art (SOTA) diffusion-based depth estimators (e.g., Marigold (Ke et al., 2024)) on NYUv2 and ScanNet, existing benchmarks often lack scene diversity, as noted in Depth Anything V2 (Yang et al., 2024b). To address this limitation, we further evaluate UniCon-Depth on the relative depth estimation benchmark DA-2K, introduced in Depth Anything V2, which comprises diverse high-resolution test images.
On DA-2K (Tab. 5), Our UniCon-Depth model is superior to other models except for Depth Anything V2 (which is used as our annotator), demonstrating that our model better generalizes to diverse scenes.
Computation efficiency analysis. We provide an analysis of our inference cost. Tab. 7 compares UniCon with ControlNet for controllable image generation. UniCon incurs additional inference overhead due to the separate computation of LoRA layers (condition LoRA and joint cross-attention LoRA) instead of fusing them into the pretrained weights. Therefore, separate LoRA layer computation incurs heavy overhead. This overhead can potentially be mitigated by saving multiple attention weights (with and without LoRA) directly in the model.
Tab. 8 compares UniCon with other depth estimation methods. Diffusion-based methods (e.g., UniCon and Marigold) generally exhibit higher latency than traditional approaches. However, UniCon can leverage techniques like Latent Consistency Models (LCM), as demonstrated by Marigold, to significantly improve inference speed.
In summary, while UniCon currently has higher inference costs, its computational efficiency can be enhanced through optimizations like LoRA weight fusion and LCM integration. We plan to explore these directions in future work.
Inpainting and editing using UniCon. UniCon supports diverse editing and inpainting behavior on both image input and condition input, as shown in Fig. 10. In addition to inpainting or editing one input conditioning on another, we can simultaneously edit and inpaint both inputs.
Empirically, we observe that our model is robust for editing tasks. However, its inpainting performance is less stable, which is a limitation of the base model (Stable Diffusion). A straightforward solution is to apply our method to an inpainting-specific diffusion model, such as SD-Inpainting.
@@
Figure 12: Qualitative results on depth estimation. We compare UniCon-Depth against Depth Anything V2 (Yang et al., 2024b), Marigold (Ke et al., 2024). Test images are from Depth Anything V2 (Yang et al., 2024b).
Figure 13: Qualitative results on conditional image generation. We compare our conditional image generation results against ControlNet (Zhang et al., 2023b), Uni-ControlNet (Zhao et al., 2024). Uni-ControlNet accepts different inputs for depth and softedge, generated by its default preprocessors (MiDaS, HED).
Interpolating between noise schedules. In Fig. 11, we show that we can interpolate the initial noise levels for input ${\mathbf{x}},{\mathbf{y}}$ to gradually change our sampling behavior in ${\mathbf{y}}$-to-${\mathbf{x}}$, rough ${\mathbf{y}}$-to-${\mathbf{x}}$, joint ${\mathbf{x}},{\mathbf{y}}$ generation, rough ${\mathbf{x}}$-to-${\mathbf{y}}$, and ${\mathbf{x}}$-to-${\mathbf{y}}$.
Qualitative results on depth estimation. We provide more qualitative results on depth estimation for our UniCon-Depth model (Fig. 12). Our method performs better than Marigold (Ke et al., 2024) on most test cases, especially for images with ambiguous structure (Fig. 12 Line 2,5) or non-realistic style (Fig. 12 Line 1,7,8). By learning the bidirectional correlation between image and depth, our model better preserves the natural image knowledge of the base model, while fine-tuning the whole model in a conditional sampling manner (as Marigold) might destroy the pretrained image prior.
Qualitative results on conditional image generation. We provide more qualitative results of our Depth, Edge, and Pose model (Fig. 13). ControlNet (Zhang et al., 2023b) sometimes generates over-saturated images (SoftEdge 2,4 in Fig. 13) or images not aligned with the condition (Pose 4 in Fig. 13), while it rarely happens to our method. We attribute it to that joint modeling stimulates our model to fully capture the bidirectional correlation between image and condition, while conditional modeling might learn to rely on minor clues in the condition image.